# positioning

### v1

@@ 5. Hodge theoretic preliminaries and a quick introduction to Zilber–Pink
We begin this section by recalling the André–Oort and Zilber–Pink conjectures and some formalism from Hodge theory. Loosely speaking, the results on (closed) geodesics are inspired by the analogy between closed geodesics and CM points [^5] A point of a Shimura variety is called CM if its associated Mumford–Tate group is a torus. and the following.
@@
It might be interesting to compare this to what happens with abelian varieties. If $A,B$ are complex abelian varieties, any morphism $f:A\to B$ has totally geodesic image. This follows from a simple Rigidity lemma, as in [32, Cor 1, page 43]. If $f$ sends a torsion point of $A$ to a torsion point of $B$, then the aforementioned corollary shows that $f$ is in fact a homomorphism. This latter statement is the analogue of 5.1 and shows that a version of Siu’s immersion problem, 1.2, holds for abelian varieties.
@@ 5.1. Hodge theory and period domains
The category of pure $\mathbb{Q}$-Hodge-structures is Tannakian and moreover semisimple when considering polarizable Hodge structures, as we frequently will. The Mumford–Tate group $\mathbf{MT}(V)\subset\mathbf{GL}(V)$ of a $\mathbb{Q}$-Hodge structure $V$ is the Tannakian group of the Tannakian subcategory $\langle V\rangle^{\otimes}$ of $\mathbb{Q}$-Hodge structures generated by $V$. Equivalently, $\mathbf{MT}(V)$ is the smallest $\mathbb{Q}$-algebraic subgroup of $\mathbf{GL}(V)$ whose base-change to $\mathbb{R}$ contains the image of $h:\mathbb{S}\to\mathbf{GL}(V_{\mathbb{R}})$. It is also the stabilizer in $\mathbf{GL}(V)$ of the Hodge tensors for $V$. As $V$ is polarized, this is a reductive algebraic group.
@@ 5.2. Typical and atypical intersections
(5.2.1) $\Phi:S^{\textnormal{an}}\longrightarrow\Gamma\backslash D,$
@@
(5.2.2) $\Phi:S^{\textnormal{an}}\longrightarrow\Gamma\backslash D\cong\Gamma_{1}\backslash D_{1}\times\cdots\times\Gamma_{r}\backslash D_{r}\times D_{\mathrm{triv}},$
@@
    A subvariety $Z$ of $S$ is said to be of positive period dimension for ${\mathbb{V}}$ if $\Phi(Z^{\textnormal{an}})$ has positive dimension, that is, $\dim_{\mathbb{C}}\Phi(Z^{\textnormal{an}})>0$.
@@
(5.2.3) $\operatorname{codim}_{\Gamma\backslash D}\Phi(Z^{\textnormal{an}})<\operatorname{codim}_{\Gamma\backslash D}\Phi(S^{\textnormal{an}})+\operatorname{codim}_{\Gamma\backslash D}\Gamma^{\prime}\backslash D^{\prime}.$
@@ 5.3. Some results on the distribution of the Hodge locus
The above statement already hides an application of 5.8, as explained in [7, Rmk. 10.2]. Indeed it is proved by first showing that $\textnormal{HL}(S,{\mathbb{V}}^{\otimes})_{\textnormal{pos}}$ in dense then invoking 5.8 to say that $\textnormal{HL}(S,{\mathbb{V}}^{\otimes})_{\textnormal{pos},\textnormal{atyp}}$ is algebraic, and therefore is exactly $\textnormal{HL}(S,{\mathbb{V}}^{\otimes})_{\textnormal{pos},\textnormal{typ}}$, which is dense. In future applications we will actually use this finer version. See also [23, Thm. 1.6 and Rmk. 1.7] for a related discussion.
@@ 6. General Hodge theoretic statement: from typical to atypical intersections
6.1. Isogenies between VHS
For an introduction to Tannakian categories, we refer for example to the article of Deligne and Milne [16]. The category of $\mathbb{Q}$VHS on a smooth quasi-projective base $S$, that we denote by $\mathbb{Q}VHS/S$, is Tannakian (we can fix a fiber functor $\omega_{s}$ corresponding to some base point $s\in S$). Given a $\mathbb{Z}VHS$ ${\mathbb{V}}$, we denote the associated $\mathbb{Q}VHS$ by ${\mathbb{V}}_{\mathbb{Q}}$.
@@ 6.1. Isogenies between VHS
Let $\mathbb{V}_{1},{\mathbb{V}}_{2}\in\mathbb{Z}VHS/S$. We say that ${\mathbb{V}}_{1}$ and ${\mathbb{V}}_{2}$ are isogenous if there is an isomorphism $\langle{\mathbb{V}}_{1,\mathbb{Q}}\rangle^{\otimes}\cong\langle{\mathbb{V}}_{2,\mathbb{Q}}\rangle^{\otimes}$ as Tannakian categories, where $\langle{\mathbb{V}}_{i,\mathbb{Q}}\rangle^{\otimes}$ denotes the smallest Tannakian subcategory of $\mathbb{Q}VHS$’s containing ${\mathbb{V}}_{i,\mathbb{Q}}$.
@@
Two complex principally polarized abelian varieties $A,B$ that are isogenous in the usual sense (either via a polarized or an unpolarized isogeny), are also isogenous in the sense of 6.1. However the converse is in general not true, for example $H^{1}(A,\mathbb{Z})$ is isogenous to $H^{1}(A\times A,\mathbb{Z})$. However, for two principally polarized $g$-dimensional abelian varieties with Mumford–Tate group $\mathbf{GSp}_{2g}$ the two notions agree, since this is the case when the Mumford–Tate group is as big as possible.
@@
such that the two projections $p_{1}:V\rightarrow\Gamma_{1}\backslash D_{1}$ and $p_{2}:V\rightarrow\Gamma_{2}\backslash D_{2}$ are finite étale surjective maps (see also [7, Def. 3.23] for more details). In this situation $V$ is a Mumford–Tate domain, that is $V=\Gamma\backslash D$, we have that $D\simeq D_{1}\simeq D_{2}$ and that $\Gamma_{1}$ and $\Gamma_{2}$ are commensurable. The following proposition translates the isogeny relation from a Tannakian statement to a more useful one on period maps.
@@
Let ${\mathbb{V}}_{1},{\mathbb{V}}_{2}$ be $\mathbb{Z}VHS$ on $S$, with associated Hodge data $({\mathbf{G}}_{i},D_{i})$ and period map $\psi_{i}$. The following are equivalent:
@@
    We have $({\mathbf{G}}_{1},D_{1})$, $({\mathbf{G}}_{2},D_{2})$ are isomorphic as Hodge datum. Using $({\mathbf{G}},D)$ for the corresponding Hodge datum and $\psi_{1}:S\to\Gamma_{1}\backslash D$, $\psi_{2}:S\to\Gamma_{2}\backslash D$ for the period maps associated with ${\mathbb{V}}_{1}$, ${\mathbb{V}}_{2}$ (respectively), then there exists a modular correspondence
@@
Let ${\mathbb{V}}$ be a $\mathbb{Z}$VHS on $S$ with generic Hodge datum $({\mathbf{G}},D)$. Then the Tannakian category generated by ${\mathbb{V}}_{\mathbb{Q}}$ is determined by $({\mathbf{G}},D)$ as any ${\mathbb{W}}$ in $\langle{\mathbb{V}}_{\mathbb{Q}}\rangle^{\otimes}$ is, by definition of a Tannakian category, given by a representation of ${\mathbf{G}}$ on the fibre $W_{\mathbb{Q}}$ of ${\mathbb{W}}_{\mathbb{Q}}$ at a point $s\in S$ and the Hodge structure on $W_{\mathbb{Q}}$ is determined by some Hodge morphism $\alpha:\mathbb{S}\rightarrow{\mathbf{G}}$ belonging to $D$. Therefore if $({\mathbf{G}}_{1},D_{1})\cong({\mathbf{G}}_{2},D_{2})$, then ${\mathbb{V}}_{1},{\mathbb{V}}_{2}$ are isogenous.
@@ 6.2. The main statement
We remark here that, in fact, some stronger variants with the weakly special Hodge locus replacing the Hodge locus can be given.
@@
(6.2.1) $\textnormal{HL}(S,{\mathbb{V}}_{1}^{\otimes})_{\textnormal{pos},\textnormal{typ}}=\textnormal{HL}(S,{\mathbb{V}}_{2}^{\otimes})_{\textnormal{pos},\textnormal{typ}}\neq\emptyset,$
@@
We will use 6.2 and the following fact: Given two principally polarized families of abelian varieties $h_{i}:\mathcal{A}_{i}\to S$, one has $\operatorname{Hom}_{S}(\mathcal{A}_{1},\mathcal{A}_{2})=\operatorname{Hom}(R_{1}{h_{1},_{*}}\mathbb{Z},R_{1}{h_{2},_{*}}\mathbb{Z})$; see [15, Cor. 4.4.15]. Indeed 6.4 implies that the two $\mathbb{Z}$VHSs ${\mathbb{V}}_{i}$ for $i\in\{1,2\}$ are isogenous in the sense of 6.1, and the above facts imply that this notion of isogeny translates precisely to the traditional one. ∎
@@
The above statement bears some similarities with work of Baker and DeMarco [4] in complex dynamics, who show that for any fixed $a,b\in\mathbb{C}$ and any integer $d\geq 0$, the set of $c\in\mathbb{C}$ for which both $a$ and $b$ are preperiodic for $z^{d}+c$ is infinite if and only if $a^{d}=b^{d}$. Also see generalizations like in the work of Yuan and Zhang [42]. Both proofs mentioned above use equidistribution but here we work with positive dimensional subvarieties. The statement without pos is deeper and would follow from 5.6.
@@
The main step in the proof is the following lemma for which we do not need to assume that the generic Mumford–Tate group of ${\mathbb{V}}_{i}$ is $\mathbb{Q}$-simple.
@@
Let $f_{i}:S\to\Gamma_{i}\backslash D_{i}$ be the period map associated with $\mathbb{V}_{i}$, $i=1,2$, and let
@@
As $\textnormal{HL}(S,{\mathbb{V}}_{i}^{\otimes})_{\textnormal{pos},\textnormal{typ}}\neq\emptyset$, by 5.9 there exists sequences of special subvarieties $\{Z_{i,\ell}\}_{\ell\in\mathbb{N}}$ in $\Gamma_{i}\backslash D_{i}$ and a Zariski dense sequence $\{W_{\ell}\}_{\ell\in\mathbb{N}}$ of components of both $\textnormal{HL}(S,{\mathbb{V}}_{1}^{\otimes})_{\textnormal{pos},\textnormal{typ}}$ and $\textnormal{HL}(S,{\mathbb{V}}_{2}^{\otimes})_{\textnormal{pos},\textnormal{typ}}$ such that
@@
for some components $f_{i}^{-1}(Z_{i,\ell})^{0}$ of $f_{i}^{-1}(Z_{i,\ell})$. The fact that $W_{\ell}$ are again typical intersections, guaranteed by 5.9, implies that $W_{\ell}$ is Hodge generic for ${\mathbb{V}}_{i}$ in $Z_{i,\ell}$ (see also the discussion in 5.5). By passing to a subsequence, we assume that $Z_{i,\ell}$ has fixed dimension $z_{i}$.
@@
$\operatorname{codim}_{\Gamma_{1}\backslash D_{1}\times\Gamma_{2}\backslash D_{2}}(f_{1}\times f_{2}(W_{\ell}^{\prime}))<\operatorname{codim}_{\Gamma_{1}\backslash D_{1}\times\Gamma_{2}\backslash D_{2}}(Z_{1,\ell}\times Z_{2,\ell})+\operatorname{codim}_{\Gamma_{1}\backslash D_{1}\times\Gamma_{2}\backslash D_{2}}(f_{1}\times f_{2}(S)).$
Summing Equation (6.2.2) for $i=1,2$ we have
(6.2.3) $d_{1}+d_{2}-w_{1}-w_{2}=d_{1}+d_{2}-z_{1}-z_{2}+d_{1}+d_{2}-s_{1}-s_{2}.$
@@
    If the monodromy of ${\mathbb{V}}_{i}$ is equal to the derived subgroup of its Mumford–Tate group ${\mathbf{G}}_{i}$, then $f_{1}\times f_{2}(S)$ is not Hodge generic in $\Gamma_{1}\backslash D_{1}\times\Gamma_{2}\backslash D_{2}$.
@@
    If the Mumford–Tate group ${\mathbf{G}}_{i}$ of ${\mathbb{V}}_{i}$ is $\mathbb{Q}$-simple, then ${\mathbf{G}}_{1}^{\textnormal{ad}}\cong{\mathbf{G}}_{2}^{\textnormal{ad}}$, $D_{1}\simeq D_{2}$. Denoting by $({\mathbf{G}},D)$ the corresponding Hodge datum, we have that $f_{1}\times f_{2}(S)$ is Hodge generic in a modular correspondence
@@
By 6.8, $S$ has a Zariski dense set of positive dimensional atypical components of the Hodge locus for the period map $f_{1}\times f_{2}$. Therefore $S$ satisfies conclusion (a) or (b) of 5.8. If conclusion (a) holds, then $f_{1}\times f_{2}(S)$ is not Hodge generic in $\Gamma_{1}\backslash D_{1}\times\Gamma_{2}\backslash D_{2}$, as desired. If the monodromy of ${\mathbb{V}}_{i}$ is equal to the derived subgroup of its Mumford–Tate group ${\mathbf{G}}_{i}$, then the projection of $f_{i}(S)$ on every factor of $\Gamma_{i}\backslash D_{i}$ is positive dimensional and Hodge generic. The same is true for the period map $f_{1}\times f_{2}$. Therefore conclusion (b) is not satisfied. This finishes the proof of the first part of the lemma.
If the Mumford–Tate group ${\mathbf{G}}_{i}$ of ${\mathbb{V}}_{i}$ is $\mathbb{Q}$-simple, then the monodromy of ${\mathbb{V}}_{i}$ is equal to the derived subgroup of its Mumford–Tate group ${\mathbf{G}}_{i}$ and by the first part, $f_{1}\times f_{2}(S)$ is not Hodge generic in $\Gamma_{1}\backslash D_{1}\times\Gamma_{2}\backslash D_{2}$. Moreover in this situation, the only strict special subvarieties of $\Gamma_{1}\backslash D_{1}\times\Gamma_{2}\backslash D_{2}$ whose projections on both factors are surjective are the modular correspondences. This forces $D_{1}\simeq D_{2}$ and ${\mathbf{G}}_{1}^{\textnormal{ad}}\cong{\mathbf{G}}_{2}^{\textnormal{ad}}$. This finishes the proof of the second part of the lemma. ∎
@@
6.4 implies 4.1 on ball quotients, which we now explain. On the smaller ball quotient $S_{n}$, which we assume to be arithmetic for simplicity, we have a $\mathbb{Z}$VHS $\mathbb{V}_{1}$ induced by any faithful linear representation of $\operatorname{GU}(1,n)$. The Hodge locus for such a VHS does not depend on the linear representation by a theorem of Deligne and it is equal to the union of all sub-Shimura varieties. With the terminology introduced in 5.4, a sub-Shimura variety is always in the typical Hodge locus. In particular
$\textnormal{HL}(S,\mathbb{V}_{1}^{\otimes})_{\textnormal{pos}}=\textnormal{HL}(S,\mathbb{V}_{1}^{\otimes})_{\textnormal{pos},\textnormal{typ}}$
and this locus is equal to the union of all totally geodesic subvarieties. Now a map $f:S_{n}\to S_{m}$ gives another VHS $\mathbb{V}_{2}$ by pulling back the natural one we have on $S_{m}$. The intersection theoretic assumptions appearing in 4.1 force $\textnormal{HL}(S,\mathbb{V}_{1}^{\otimes})_{\textnormal{pos},\textnormal{typ}}=\textnormal{HL}(S,\mathbb{V}_{2}^{\otimes})_{\textnormal{pos},\textnormal{typ}}$. Therefore we conclude by applying 6.4.
7. Proofs of 1.11 and 1.14
In this section we prove the second part of 1.11 and then 1.14 following a similar argument. Let $\Gamma\subset\operatorname{PU}(1,n)$ for $n>1$ be a lattice of the first kind, which we can assume to be neat. Consider the smooth quasi-projective variety $S_{\Gamma}$ and let
@@
is the adjoint representation over some local field $k$. Assuming that $\rho$ is geodesically rich in the sense of 1.6, we would like to show that $\rho$ extends. The proof is a generalization of the arguments from [8, Sec. 4 and 5], improved by combining them with Section 6.
Proof of 1.11 (2).

### v3

@@ 5. Hodge theoretic preliminaries and a quick introduction to Zilber–Pink
We begin this section by recalling the André–Oort and Zilber–Pink conjectures and some formalism from Hodge theory. Loosely speaking, our results on (closed) geodesics are inspired by the analogy between closed geodesics and CM points [^5] A point of a Shimura variety is called CM if its associated Mumford–Tate group is a torus. and the following.
@@
It may be interesting to compare this to what happens with abelian varieties. If $A,B$ are complex abelian varieties, any morphism $f:A\to B$ has totally geodesic image. This follows from a simple Rigidity lemma as in [35, Cor 1, page 43]. If $f$ sends a torsion point of $A$ to a torsion point of $B$, then the aforementioned corollary shows that $f$ is in fact a homomorphism. This latter statement is the analogue of 5.1 and shows that a version of Siu’s immersion problem (1.2) holds for abelian varieties.
@@ 5.1. Hodge theory and period domains
The category of pure $\mathbb{Q}$-Hodge-structures is Tannakian and moreover semisimple when considering polarizable Hodge structures, as we frequently will. The Mumford–Tate group $\mathbf{MT}(V)\subset\mathbf{GL}(V)$ of a $\mathbb{Q}$-Hodge structure $V$ is the Tannakian group of the Tannakian subcategory $\langle V\rangle^{\otimes}$ of $\mathbb{Q}$-Hodge structures generated by $V$. Equivalently, $\mathbf{MT}(V)$ is the smallest $\mathbb{Q}$-algebraic subgroup of $\mathbf{GL}(V)$ whose base-change to $\mathbb{R}$ contains the image of $h:\mathbb{S}\to\mathbf{GL}(V_{\mathbb{R}})$. It is also the stabilizer in $\mathbf{GL}(V)$ of the Hodge tensors for $V$. When $V$ is polarized, this is a reductive algebraic group.
@@ 5.2. Typical and atypical intersections
(5.1) $\Phi:S^{\textnormal{an}}\longrightarrow\Gamma\backslash D,$
@@
(5.2) $\Phi:S^{\textnormal{an}}\longrightarrow\Gamma\backslash D\cong\Gamma_{1}\backslash D_{1}\times\cdots\times\Gamma_{r}\backslash D_{r}\times D_{\mathrm{triv}},$
@@
    A subvariety $Z$ of $S$ is said to be of positive period dimension for ${\mathbb{V}}$ if $\Phi(Z^{\textnormal{an}})$ has positive dimension, that is, if $\dim_{\mathbb{C}}\Phi(Z^{\textnormal{an}})>0$.
@@
(5.3) $\operatorname{codim}_{\Gamma\backslash D}\Phi(Z^{\textnormal{an}})<\operatorname{codim}_{\Gamma\backslash D}\Phi(S^{\textnormal{an}})+\operatorname{codim}_{\Gamma\backslash D}\Gamma^{\prime}\backslash D^{\prime}.$
@@ 5.3. Some results on the distribution of the Hodge locus
We refer also to [9, Thm. 7.1] for an effective proof of a more general statement (proved using differential geometry rather than $o$-minimality).
@@
The above statement already hides an application of 5.8, as explained in [7, Rmk. 10.2]. Indeed it is proved by first showing that $\textnormal{HL}(S,{\mathbb{V}}^{\otimes})_{\textnormal{pos}}$ in dense then invoking 5.8 to say that $\textnormal{HL}(S,{\mathbb{V}}^{\otimes})_{\textnormal{pos},\textnormal{atyp}}$ is algebraic, and therefore is exactly $\textnormal{HL}(S,{\mathbb{V}}^{\otimes})_{\textnormal{pos},\textnormal{typ}}$, which is dense. In future applications we will actually use this finer version. See also [24, Thm. 1.6 and Rmk. 1.7] for a related discussion.
@@ 6. General Hodge theoretic statement: from typical to atypical intersections
6.1. Isogenies between VHSs
For an introduction to Tannakian categories, we refer for example to the article of Deligne and Milne [17]. The category of $\mathbb{Q}$VHS on a smooth quasi-projective base $S$, which we denote by $\mathbb{Q}VHS/S$, is Tannakian (we can fix a fiber functor $\omega_{s}$ corresponding to some base point $s\in S$). Given a $\mathbb{Z}VHS$ ${\mathbb{V}}$, we denote the associated $\mathbb{Q}VHS$ by ${\mathbb{V}}_{\mathbb{Q}}$.
@@ 6.1. Isogenies between VHSs
Let $\mathbb{V}_{1},{\mathbb{V}}_{2}\in\mathbb{Z}VHS/S$. We say that ${\mathbb{V}}_{1}$ and ${\mathbb{V}}_{2}$ are isogenous if there is an equivalence of tensor categories $\langle{\mathbb{V}}_{1,\mathbb{Q}}\rangle^{\otimes}\cong\langle{\mathbb{V}}_{2,\mathbb{Q}}\rangle^{\otimes}$, where $\langle{\mathbb{V}}_{i,\mathbb{Q}}\rangle^{\otimes}$ denotes the smallest Tannakian subcategory of $\mathbb{Q}VHS$’s containing ${\mathbb{V}}_{i,\mathbb{Q}}$.
Of course the Tannakian categories $\langle{\mathbb{V}}_{i,\mathbb{Q}}\rangle^{\otimes}$ appearing above are equivalent (as tensor categories) to the category of finite dimensional representations of their generic Mumford-Tate group. (The equivalence is realized by the functor of $\otimes$-automorphisms of the fiber functor). It can happen that two VHS have isomorphic Mumford-Tate groups, but the isomorphism does not induce an equivalence of tenor categories. C.f. Def. 1.10 and Thm. 2.11 in the article of Deligne and Milne[17].
@@
If two complex principally polarized abelian varieties $A,B$ that are isogenous in the usual sense (either via a polarized or an unpolarized isogeny), then the Hodge structures $H^{1}(A,\mathbb{Z}),H^{1}(B,\mathbb{Z})$ also isogenous in the sense of 6.1 (here we take as base $S$ the spectrum of $\mathbb{C}$). However the converse is in general not true, for example $H^{1}(A,\mathbb{Z})$ is isogenous to $H^{1}(A\times A,\mathbb{Z})$. However, for two principally polarized $g$-dimensional abelian varieties with Mumford–Tate group $\mathbf{GSp}_{2g}$ the two notions agree, since this is the case when the Mumford–Tate group is as big as possible.
@@
such that the two projections $p_{1}:V\rightarrow\Gamma_{1}\backslash D_{1}$ and $p_{2}:V\rightarrow\Gamma_{2}\backslash D_{2}$ are finite étale surjective maps (see also [7, Def. 3.23] for more details). In this situation $V$ is a Mumford–Tate domain, that is $V=\Gamma\backslash D$, we have that $D\simeq D_{1}\simeq D_{2}$ and that $\Gamma_{1}$ is commensurable with $\Gamma_{2}$. The following proposition translates the isogeny relation from a Tannakian statement to a more useful one on period maps.
@@
Let ${\mathbb{V}}_{1},{\mathbb{V}}_{2}$ be $\mathbb{Z}VHS$ on $S$ with associated Hodge data $({\mathbf{G}}_{i},D_{i})$ and period map $\psi_{i}$. The following are equivalent:
@@
    The Hodge data $({\mathbf{G}}_{1},D_{1})$, $({\mathbf{G}}_{2},D_{2})$ are isomorphic. Using $({\mathbf{G}},D)$ for the corresponding Hodge datum and $\psi_{1}:S\to\Gamma_{1}\backslash D$, $\psi_{2}:S\to\Gamma_{2}\backslash D$ for the period maps associated with ${\mathbb{V}}_{1}$, ${\mathbb{V}}_{2}$ (respectively), then there exists a modular correspondence
@@
Let ${\mathbb{V}}$ be a $\mathbb{Z}$VHS on $S$ with generic Hodge datum $({\mathbf{G}},D)$. Then the Tannakian category generated by ${\mathbb{V}}_{\mathbb{Q}}$ is determined by $({\mathbf{G}},D)$, as any ${\mathbb{W}}$ in $\langle{\mathbb{V}}_{\mathbb{Q}}\rangle^{\otimes}$ is, by definition of a Tannakian category, given by a representation of ${\mathbf{G}}$ on the fibre $W_{\mathbb{Q}}$ of ${\mathbb{W}}_{\mathbb{Q}}$ at a point $s\in S$ and the Hodge structure on $W_{\mathbb{Q}}$ is determined by some Hodge morphism $\alpha:\mathbb{S}\rightarrow{\mathbf{G}}$ belonging to $D$. Therefore if $({\mathbf{G}}_{1},D_{1})\cong({\mathbf{G}}_{2},D_{2})$, then ${\mathbb{V}}_{1},{\mathbb{V}}_{2}$ are isogenous.
@@ 6.2. The main statement
Some stronger variants with the weakly special Hodge locus replacing the Hodge locus can be given. Also, the same proof works if there exits a Zariski dense subset of components of $\textnormal{HL}(S,\mathbb{V}_{1}^{\otimes})_{\textnormal{pos},\textnormal{typ}}$ that is contained in $\textnormal{HL}(S,\mathbb{V}_{2}^{\otimes})_{\textnormal{pos},\textnormal{typ}}$.
@@
(6.1) $\textnormal{HL}(S,{\mathbb{V}}_{1}^{\otimes})_{\textnormal{pos},\textnormal{typ}}=\textnormal{HL}(S,{\mathbb{V}}_{2}^{\otimes})_{\textnormal{pos},\textnormal{typ}}\neq\emptyset,$
@@
We will use 6.2 and the following fact: Given two principally polarized families of abelian varieties $h_{i}:\mathcal{A}_{i}\to S$, one has $\operatorname{Hom}_{S}(\mathcal{A}_{1},\mathcal{A}_{2})=\operatorname{Hom}(R_{1}{h_{1},_{*}}\mathbb{Z},R_{1}{h_{2},_{*}}\mathbb{Z})$, where the second $\operatorname{Hom}$ is in the category of $\mathbb{Z}$VHS; see [16, Cor. 4.4.15]. Indeed 6.4 implies that the two $\mathbb{Z}$VHSs ${\mathbb{V}}_{i}$ for $i\in\{1,2\}$ are isogenous in the sense of 6.1, and the above facts imply that this notion of isogeny translates precisely to the traditional one. ∎
@@
The above statement bears some similarities with work of Baker and DeMarco [4] in complex dynamics, which shows that for any fixed $a,b\in\mathbb{C}$ and any integer $d\geq 0$, the set of $c\in\mathbb{C}$ for which both $a$ and $b$ are preperiodic for $z^{d}+c$ is infinite if and only if $a^{d}=b^{d}$. Also see generalizations like in the work of Yuan and Zhang [45]. Both proofs mentioned above use equidistribution but here we work with positive dimensional subvarieties. The statement without pos is deeper and would follow from 5.6.
@@
The main step in the proof is the following lemma, where we do not need to assume that the generic Mumford–Tate group of ${\mathbb{V}}_{i}$ is $\mathbb{Q}$-simple.
@@
Assuming the hypothesis of 6.4, let $f_{i}:S\to\Gamma_{i}\backslash D_{i}$ be the period map associated with $\mathbb{V}_{i}$, $i=1,2$, and let
@@
As $\textnormal{HL}(S,{\mathbb{V}}_{i}^{\otimes})_{\textnormal{pos},\textnormal{typ}}\neq\emptyset$, by 5.9 there exists sequences of special subvarieties $\{Z_{i,\ell}\}_{\ell\in\mathbb{N}}$ in $\Gamma_{i}\backslash D_{i}$ and a sequence $\{W_{\ell}\}_{\ell\in\mathbb{N}}$ of components of both $\textnormal{HL}(S,{\mathbb{V}}_{1}^{\otimes})_{\textnormal{pos},\textnormal{typ}}$ and $\textnormal{HL}(S,{\mathbb{V}}_{2}^{\otimes})_{\textnormal{pos},\textnormal{typ}}$, such that
@@
for some components $f_{i}^{-1}(Z_{i,\ell})^{0}$ of $f_{i}^{-1}(Z_{i,\ell})$ and for which $\cup_{\ell}W_{\ell}$ is Zariski dense in $S$. The fact that the $W_{\ell}$ are again typical intersections, guaranteed by 5.9, implies that $W_{\ell}$ is Hodge generic for ${\mathbb{V}}_{i}$ in $Z_{i,\ell}$ (see also the discussion in 5.5). By passing to a subsequence, we assume that $Z_{i,\ell}$ has fixed dimension $z_{i}$.
@@
$\operatorname{codim}_{\Gamma_{1}\backslash D_{1}\times\Gamma_{2}\backslash D_{2}}((f_{1}\times f_{2})(W_{\ell}^{\prime}))<\operatorname{codim}_{\Gamma_{1}\backslash D_{1}\times\Gamma_{2}\backslash D_{2}}(Z_{1,\ell}\times Z_{2,\ell})+\operatorname{codim}_{\Gamma_{1}\backslash D_{1}\times\Gamma_{2}\backslash D_{2}}((f_{1}\times f_{2})(S)).$
Summing Equation (6.2) for $i=1,2$ we have
(6.3) $d_{1}+d_{2}-w_{1}-w_{2}=d_{1}+d_{2}-z_{1}-z_{2}+d_{1}+d_{2}-s_{1}-s_{2}.$
@@
    If the monodromy of ${\mathbb{V}}_{i}$ is equal to the derived subgroup of its Mumford–Tate group ${\mathbf{G}}_{i}$, then $(f_{1}\times f_{2})(S)$ is not Hodge generic in $\Gamma_{1}\backslash D_{1}\times\Gamma_{2}\backslash D_{2}$.
@@
    If the Mumford–Tate group ${\mathbf{G}}_{i}$ of ${\mathbb{V}}_{i}$ is $\mathbb{Q}$-simple, then ${\mathbf{G}}_{1}^{\textnormal{ad}}\cong{\mathbf{G}}_{2}^{\textnormal{ad}}$, $D_{1}\simeq D_{2}$. Denoting the corresponding Hodge datum by $({\mathbf{G}},D)$, we have that $(f_{1}\times f_{2})(S)$ is Hodge generic in a modular correspondence
@@
By 6.8, $S$ has a Zariski dense set of positive dimensional atypical components of the Hodge locus for the period map $f_{1}\times f_{2}$. Therefore $S$ satisfies conclusion (a) or (b) of 5.8. If conclusion (a) holds, then $(f_{1}\times f_{2})(S)$ is not Hodge generic in $\Gamma_{1}\backslash D_{1}\times\Gamma_{2}\backslash D_{2}$, as desired. If the monodromy of ${\mathbb{V}}_{i}$ is equal to the derived subgroup of its Mumford–Tate group ${\mathbf{G}}_{i}$, then the projection of $f_{i}(S)$ on every factor of $\Gamma_{i}\backslash D_{i}$ is positive dimensional and Hodge generic. The same is true for the period map $f_{1}\times f_{2}$. Therefore conclusion (b) is not satisfied. This finishes the proof of the first part of the lemma.
If the Mumford–Tate group ${\mathbf{G}}_{i}$ of ${\mathbb{V}}_{i}$ is $\mathbb{Q}$-simple, then the monodromy of ${\mathbb{V}}_{i}$ is equal to the derived subgroup of its Mumford–Tate group ${\mathbf{G}}_{i}$ and by the first part, $(f_{1}\times f_{2})(S)$ is not Hodge generic in $\Gamma_{1}\backslash D_{1}\times\Gamma_{2}\backslash D_{2}$. Moreover in this situation, the only strict special subvarieties of $\Gamma_{1}\backslash D_{1}\times\Gamma_{2}\backslash D_{2}$ whose projections on both factors are surjective are the modular correspondences. This forces $D_{1}\simeq D_{2}$ and ${\mathbf{G}}_{1}^{\textnormal{ad}}\cong{\mathbf{G}}_{2}^{\textnormal{ad}}$. This finishes the proof of the second part of the lemma. ∎
@@
6.4 (or, more precisely, a slight generalization of it) implies 4.1 for ball quotients, which we now explain. On the smaller ball quotient $S_{n}$, which we assume to be arithmetic for simplicity, we have a $\mathbb{Z}$VHS $\mathbb{V}_{1}$ induced by any faithful linear representation of $\operatorname{GU}(1,n)$. The Hodge locus for such a VHS does not depend on the linear representation by a theorem of Deligne and it is equal to the union of all sub-Shimura varieties. With the terminology introduced in 5.4, a sub-Shimura variety is always in the typical Hodge locus
$\textnormal{HL}(S_{n},\mathbb{V}_{1}^{\otimes})_{\textnormal{pos}}=\textnormal{HL}(S_{n},\mathbb{V}_{1}^{\otimes})_{\textnormal{pos},\textnormal{typ}}$
and this locus is equal to the union of all totally geodesic subvarieties. Now a map $f:S_{n}\to S_{m}$ gives another VHS $\mathbb{V}_{2}$ by pulling back the natural one on $S_{m}$. Once this translation from immersions to Hodge theory is made, the proof of 4.1 is just a special case of the one given in 6.4.
7. Proofs of 1.13 and 1.17
In this section we prove the second part of 1.13 and then 1.17 following a similar argument. Let $\Gamma\subset\operatorname{PU}(1,n)$ for $n>1$ be a lattice of the simplest kind, which we can assume to be neat (i.e., no element of $\Gamma$ has a nontrivial root of unity as an eigenvalue under the adjoint representation). Consider the smooth quasi-projective variety $S_{\Gamma}$ and let
@@
is the adjoint representation over some local field $k$. Assuming that $\rho$ is geodesically rich in the sense of 1.8, we would like to show that $\rho$ extends. The proof is a generalization of the arguments from [8, Sec. 4 and 5], improved by combining them with Section 6.
Proof of 1.13 (2).